# Quantum Langevin Dynamics for Optimization

### v2

@@ 6 Discussion and Conclusions
In conclusion, we demonstrate the possibility of solving continuous optimization problems through the application of a specific class of open dissipative quantum dynamics. We have theoretical guarantee of the convergence of QLD; meanwhile, we verify our intuition and provide a feasible cooling strategy through numerical experiments. In particular, although in some landscapes other algorithms have already performed well, by controlling the parameters of QLD, it can outperform all other algorithms in some other landscapes.
@@
In particular, due to the complex formula of the quantum Langevin dynamics, we only provide the convergence of QLD under the approximations of $\hbar\Omega\ll kT$ and $T\to 0$. Since the limit $\hbar\to 0$ clearly gives a transition to the classical Langevin equation, where all commutators vanish in this limit, in a sense, we might just prove the convergence of QLD in the classical limit, and use its quantum effect in the process of convergence. Whether the convergence still exists without these approximations remains further discussion with more meticulous mathematical and physical analysis. Numerically, since the simulation in the one-dimensional landscape alone has consumed a considerable amount of our computational resources, numerical experiments in high dimension requires more efficient numerical methods and more powerful classical computers. This also demonstrates the necessity of analog (or digital) quantum computers.
@@ Acknowledgements
We thank Zhiyan Ding, Jiaqi Leng, Lin Lin, Hao Wu, Yukai Wu, Shuohan Zhang, and Zhennan Zhou for helpful discussions. Z. C. and Y. L. acknowledge research funding from Tsien Excellence in Engineering Program of Tsinghua University and Tsinghua University Initiative Scientific Research Program. T. L. was supported by the National Natural Science Foundation of China (Grant Numbers 62372006 and 92365117), and a startup fund from Peking University.
@@ A.1 Implementation details of numerical experiments for QLD
[…]
@@ A.3 Analysis of numerical results
Appendix B Comparison of query complexity for different algorithms
In this appendix, we delve into the comparison of query complexity among various algorithms, i.e., the classical query complexity to local information and the quantum query complexity to the evaluation oracle [59, 16]. We assume that the quantum evaluation oracle $O_{V}$ allows us to query the function value of $V$ in a coherent way
B.1 Quantum query complexity for QLD
Theorem 4.
Given a Lindblad equation with form Eq. (14),
$O\left(t(n^{2}+\|V\|_{\max})\frac{\log(t(n^{2}+\|V\|_{\max})/\epsilon)}{\log\log(t(n^{2}+\|V\|_{\max})/\epsilon)}\right)$
Proof.
By Definition 3, it is not difficult to give an $(\alpha_{E_{k}},a,\xi_{E_{k}})$-block-encoded matrix $U_{E_{k}}$ of $E_{k}$ and an $(\alpha_{V},a,\xi_{V})$-block-encoded matrix $U_{V}$ of $V$, where $\alpha_{E_{k}}$ is $O(\frac{1}{\Delta x^{2}})=O(n^{2})$ and $\alpha_{V}$ is $O(\|V\|_{\max})$. Then, Proposition 3 derives a $(\beta,a+b,\xi_{E_{k}}+\xi_{V})$-block-encoded matrix $\tilde{W}_{H}$ of $H$ with $\beta\geq|\alpha_{E_{k}}|+|\alpha_{V}|$. In addition, Proposition 3 also tells one query to $\tilde{W}_{H}$ can be implemented by one query to $U_{E_{k}}$ and one query to $U_{V}$.
Now, we have gotten a $(\beta,a+b,\xi_{E_{k}}+\xi_{V})$-block-encoded matrix $\tilde{W}_{H}$ of $H$. In our Lindblad equation Eq. (14), the unique $L$ is given by $L=\mu x+\mathrm{i}\nu p$. Presenting $L$ in the matrix form by finite difference method, we have $L=\mu\cdot\text{diag}(x_{1},\ldots,x_{n})+\nu\hbar\cdot\frac{\mathbf{D}}{\Delta x}$, where $\mathbf{D}$ is the differential matrix associated to the central difference method. Then, we can easily give an $(\alpha_{L},a,\xi_{L})$-block-encoded matrix $U_{L}$ of $L$, where $\alpha_{L}$ is at most $O(\frac{1}{\Delta x})=O(n)$.
Using Proposition 4, $\|\mathcal{L}\|_{\text{be}}=\beta+\alpha_{L}^{2}$, the smallest $\beta$ we can achieve is $|\alpha_{E_{k}}|+|\alpha_{V}|=O(n^{2})+O(\|V\|_{\max})$, which gives $\|\mathcal{L}\|_{\text{be}}=O(n^{2}+\|V\|_{\max})$. So, $O\left(t(n^{2}+\|V\|_{\max})\frac{\log(t(n^{2}+\|V\|_{\max})/\epsilon)}{\log\log(t(n^{2}+\|V\|_{\max})/\epsilon)}\right)$ queries to $\tilde{W}_{H}$ and $U_{L}$ are needed to simulate $e^{\mathcal{L}t}$ with error at most $\epsilon$. Recall that each query to $\tilde{W}_{H}$ can be implemented by one query to $U_{E_{k}}$ and one query to $U_{V}$, and one query to $U_{V}$ can be applied by one query to oracle $O_{V}$. So, $O\left(t(n^{2}+\|V\|_{\max})\frac{\log(t(n^{2}+\|V\|_{\max})/\epsilon)}{\log\log(t(n^{2}+\|V\|_{\max})/\epsilon)}\right)$ queries to $O_{V}$ are needed to simulate $e^{\mathcal{L}t}$ with error at most $\epsilon$. ∎
Definition 3 (Block-encoding, Definition 43 of [34]).
Definition 4 (State preparation pair, Definition 51 of [34]).
[…]
B.2 Query complexity of other algorithms
B.2.1 Quantum query complexity for QHD
Because QHD simulates a time-dependent Hamiltonian system, its quantum query complexity can be inferred from the complexity involved in the simulation of the time-dependent Schrödinger equation. Here, we mainly follow Ref. [15, 54] and use quantum simulation in the real space to calculate the quantum query complexity for QHD in terms of the quantum oracle $O_{V}$ (Eq. (116)).
B.2.2 Quantum query complexity for QAA
B.2.3 Classical query complexity for SGD
B.2.4 Classical query complexity for NAGD
B.3 Comparison of query complexity for different algorithms

### v3

@@ 7 Discussion and Conclusions
In conclusion, we demonstrate the possibility of solving continuous optimization problems through the application of a specific class of open dissipative quantum dynamics. We have theoretical guarantee of the convergence of QLD; meanwhile, we verify our intuition and provide a feasible cooling strategy through numerical experiments. In particular, although in some landscapes other algorithms have already performed well, by controlling the parameters of QLD, it can outperform all other algorithms in other landscapes.
@@
In particular, due to the complex formula of the quantum Langevin dynamics, we only provide the convergence of QLD under the approximations of $\hbar\Omega\ll kT$ and $T\to 0$. Since the limit $\hbar\to 0$ clearly gives a transition to the semi-classical Langevin equation, where all commutators vanish in this limit, in a sense, we might just prove the convergence of QLD in the semi-classical limit, and use its quantum effect in the process of convergence. Whether the convergence still exists without these approximations remains further discussion with more meticulous mathematical and physical analysis. Numerically, since the simulation in the one-dimensional landscape alone has consumed a considerable amount of our computational resources, numerical experiments in high dimension requires more efficient numerical methods and more powerful classical computers. This also demonstrates the necessity of analog (or digital) quantum computers.
Moreover, we only consider the jump operator as a linear combination of the position and momentum operator in this work, which follows from Ref. [28]. However, the efficacy of alternative jump operators and the potential enhancement of optimizing performance through their adoption remain open questions. In conclusion, many intriguing questions remain for further discussion, and we believe that our work will inspire more investigation into optimization using open quantum systems.
@@ Acknowledgements
We thank Zhiyan Ding, Jiaqi Leng, Lin Lin, Hao Wu, Yukai Wu, Shuohan Zhang, and Zhennan Zhou for helpful discussions.
Declarations
Funding. Z. C. and Y. L. acknowledge research funding from Tsien Excellence in Engineering Program of Tsinghua University and Tsinghua University Initiative Scientific Research Program. T. L. was supported by the National Natural Science Foundation of China (Grant Numbers 62372006 and 92365117), and a startup fund from Peking University.
Financial interests. The authors declare that they have no financial interests.
Data availability. The data that support the findings of this study are available from the corresponding author, upon reasonable request.
@@ A.1 Implementation details of numerical experiments for QLD
Nonetheless, a crucial challenge persists: determining the values of each parameter in Eq. (10). The system’s behavior is highly sensitive to the quantities chosen for all parameters in Eq. (10). Refs. [28, 29, 30] study the dynamics of vibrational relaxation of a Morse oscillator, mimic the O2-Pt(111) bond, induced by equilibrium electrons. We have nondimensionalized this real physical system to obtain the baseline values for all parameters.